Given f(x) = (Ax+5)/(6x-2) If f(1) = 4, find A
Answer by user_dude2008(1861)

f(x) = (Ax+5)/(6x-2)

4 = (A(1)+5)/(6(1)-2)

4 = (A+5)/(6-2)

4 = (A+5)/4

4*4 = A+5

16 = A+5

16-5 = A

11 = A

A = 11 <------------------ Answer